# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
The superconducting diode effect (SDE) refers to the non-reciprocal nature of the critical current (maximum current that a superconductor can withstand before turning into a normal metal) of a superconducting device. Here, we investigate SDE in helical superconductors with broken inversion and time-reversal symmetry, focusing on a prototypical Rashba nanowire device proximitized by an $s$-wave superconductor and subjected to external magnetic fields. Using a self-consistent Bogoliubov-de Gennes mean-field formalism, we analyze the interplay between linear and higher-order spin-orbit coupling (SOC), bulk supercurrents, and external magnetic fields. Our results demonstrate that Rashba nanowires with only linear SOC can achieve incredibly large diode efficiencies of $\gtrsim 45\%$ through the interplay of longitudinal and transverse magnetic fields. Notably, higher-order SOC introduces qualitatively different behavior, enabling finite diode efficiency even in the absence of a longitudinal Zeeman field due to inherent energy dispersion asymmetry. We present a comprehensive phase diagram of the device elucidating the emergent Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) superconducting state and demonstrate that proximitized Rashba nanowires offer a versatile, practical platform for SDE, with potential realizations in existing material systems. These results provide crucial insights for optimizing SDE in nanoscale superconducting devices, paving the way for next-generation dissipationless quantum electronics.
@@ I Introduction
The inversion symmetry breaking required for critical current non-reciprocity can be achieved through extrinsic mechanisms (e.g., asymmetric geometry or artificial superlattice potential Sundaresh et al. (2023); Burlakov et al. (2014); Narita and Ono (2024)) or intrinsic mechanisms like Rashba spin-orbit coupling (SOC) Ando et al. (2020); Yuan and Fu (2022); Ilić and Bergeret (2022); Halterman et al. (2022); Daido and Yanase (2022); Karabassov et al. (2022); Legg et al. (2022). Correspondingly, SDE can be broadly classified into two types: extrinsic and intrinsic. Although predicted decades ago Levitov et al. (1985); Edelstein (1996), the mechanism of SDE has only recently been clearly understood, leading to extensive investigations into its potential for dissipationless electronic circuits Yuan and Fu (2022); Daido et al. (2022); Ilić and Bergeret (2022); He et al. (2022); Pal et al. (2022); de Picoli et al. (2023); Banerjee and Scheurer (2024); Nadeem et al. (2023). Recent theoretical studies Yuan and Fu (2022); Daido et al. (2022); Ilić and Bergeret (2022); He et al. (2022), using phenomenological Ginzburg-Landau theory and mean-field analysis of the two-dimensional Rashba-Zeeman Hubbard model, have shown that intrinsic SDE can arise in systems with broken inversion and time-reversal symmetry. In the helical superconducting ground state, this occurs due to asymmetry in the FFLO Cooper pair momentum, resulting in Kramers non-degenerate $s$-wave Cooper pairs Daido et al. (2022); Yuan and Fu (2022); Nadeem et al. (2023).
@@
A Rashba nanowire device proximitized to an s-wave superconductor and subjected to external magnetic fields serves as a prototypical platform for realizing helical superconductivity Oreg et al. (2010); Sau et al. (2010); Lutchyn et al. (2010); Stanescu et al. (2011); Legg et al. (2022); de Picoli et al. (2023). While previous phenomenological/ non-self-consistent studies Legg et al. (2022) demonstrated the potential for superconducting diode effect (SDE) in such systems, they reported low diode efficiencies ($\sim 2\%$). However, to accurately determine the diode efficiency in the true superconducting ground state across the parameter space, it is crucial to systematically investigate the superconducting properties self-consistently, accounting for variations in the superconducting order parameter, as demonstrated in recent studies Daido et al. (2022); Ilić and Bergeret (2022); de Picoli et al. (2023). In this paper, we address this gap through a comprehensive, self-consistent mean-field investigation of the interplay between various Rashba SOC terms and external magnetic fields in determining emergent FFLO pairing and highly efficient superconducting diodes in a Rashba nanowire device. We show that even with only linear Rashba SOC we can achieve a very large diode efficiency ($\gtrsim 45\%$) by suitable tuning the parameters in this system. We extend beyond linear Rashba SOC to include higher-order (cubic) Rashba SOC terms, which are not only naturally present in semiconducting nanowires Campos et al. (2018); Kammermeier et al. (2018); Sonehara et al. (2022) but also can dominate the normal state properties Vajna et al. (2012); Moriya et al. (2014); Nakamura et al. (2012); Michiardi et al. (2015); Shanavas (2016); Liu et al. (2018); Peng et al. (2023); Amundsen et al. (2024). Notably, we demonstrate that higher-order Rashba SOC can induce finite diode efficiency ($\gtrsim 15\%$) even without longitudinal magnetic fields. Our study establishes the Rashba nanowire device as a promising platform for highly efficient superconducting diodes, crucial for advancing nanoscale superconducting electronics.
@@ II Model and Self-consistent mean-field formalism
We consider a single channel one-dimensional semiconducting nanowire with both linear and higher order Rashba SOC Campos et al. (2018); Kammermeier et al. (2018); Sonehara et al. (2022) and in presence of externally applied magnetic fields leading to the Rashba-Zeeman nanowire Hamiltonian (Alicea, 2012)
@@
$\!\!\hat{h}(z)=\frac{p_{z}^{2}}{2m}-\mu+\left(\frac{\alpha}{\hbar}p_{z}+\frac{\delta}{\hbar^{3}}p_{z}^{3}\right)\sigma_{x}+\frac{\gamma}{\hbar^{3}}p_{z}^{3}\sigma_{y}+\mathbf{B}.\boldsymbol{\sigma}$ (1)
Here, the spin degree of freedom is labeled by $s,s^{\prime}\in\{\uparrow,\downarrow\}$. The nanowire is placed along the $z$-direction (as shown in Fig. 1(a)) and $\psi_{s}^{\dagger}(z)$ operator creates a fermion at the position $z$ with spin $s$. $\boldsymbol{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of the Pauli matrices in the spin space and $\mu$ is the chemical potential. $p_{z}=\frac{\hbar}{i}\partial_{z}$ is the momentum operator in the direction of current flow ($z$-direction) within the nanowire. The strength of the usual linear Rashba SOC is denoted by $\alpha$ and the cubic Rashba SOC of strengths are $\delta$ and $\gamma$. The presence of Rashba SOC breaks the inversion symmetry of the system. The externally applied magnetic fields lead to the Zeeman term characterized by $\mathbf{B}=(B_{x},B_{y},B_{z})$ that breaks time-reversal symmetry. We assume the radius of the nanowire to be small such that there is only one occupied mode, i.e.,  single channel. This nanowire Hamiltonian [Eq. (1)] in momentum space can be written as
@@
$\hat{h}(k)=\xi_{k}+\left[\left(\alpha k+\delta k^{3}\right)\sigma_{x}+\gamma k^{3}\sigma_{y}\right]+\mathbf{B}.\boldsymbol{\sigma}$ (2)
@@
where, $\pm$ label the two helicty bands. Note that, the Zeeman field $B_{z}$ opens a Zeeman gap while $B_{x}$ creates asymmetry between the two bands as schematically shown in Fig. 1(b). On the other hand, $B_{y}$ opens up a gap as well as creates band asymmetry (not shown). Throughout the majority of this paper, we will discuss the results with $B_{y}=0$ for simplicity and will separately discuss the effects of nonzero $B_{y}$ for completeness.
We seek to elucidate the nature of the superconducting ground state in the Rashba nanowire in the presence of $s$-wave superconducting correlations that can be induced via the proximity effect in presence of a bulk $s$-wave superconductor, as well as supercurrents. Such a device is schematically shown in Fig. 1(a). In general, the presence of supercurrents can lead to a superconducting state with finite momentum Cooper pairs termed as the FFLO state. To establish that, we consider an attractive Hubbard type interaction in the bulk $s$-wave superconductor described by the Hamiltonian
@@
Figure 2: FFLO ground state with linear SOC: (a) The self-consistent superconducting gap $\Delta(q)$ is shown as a function of the Cooper-pair momentum $q~{}$(nm-1) for $B_{x}/\Delta_{0}=0.3$ with only linear Rashba SOC ($\alpha=1$ eV-Å). $\Delta_{0}$ is the BCS gap in the absence of any external magnetic field ($\mathbf{B}=0$). (b) The Cooper pair momentum of the FFLO ground state $q_{0}$ as a function of $B_{x}$. The presence of the FFLO ground state ($q_{0}\neq 0$) is established only when both $B_{x}$ and $B_{z}$ are finite. Other model parameters are chosen as: $(B_{y},\mu/\Delta_{0},U,\beta^{-1})$=(00, $0.5$, $25.55$ meV, $0.1$ meV).
Figure 3: FFLO ground state with both linear and higher order SOC: (a) The self-consistent superconducting gap $\Delta(q)$ is depicted as a function of Cooper-pair momentum $q~{}$ for $B_{z}=0$ with both the linear Rashba SOC term ($\alpha=1$ eV-Å) and the cubic Rashba SOC terms ($\delta=\gamma=100$ eV-Å3). (b) The Cooper pair momentum of the FFLO ground state $q_{0}~{}$ as a function of $B_{x}$ is shown. Note that the FFLO state is stable even when $B_{z}=0$ in the presence of a finite $B_{x}$. Other system parameters are chosen as: $(B_{y},\mu/\Delta_{0},U,\beta^{-1})$=(00, $0.5$, $16.45$ meV, $0.1$ meV).
@@
Figure 4: Superconducting diode effect with linear SOC: (a) The supercurrent density ($J$) is shown as a function $q~{}$(nm-1) considering solely linear Rashba SOC and $B_{x}/\Delta_{0}=0.3$. The nonreciprocity in the supercurrent, $J(q)\neq-J(-q)$, is achieved in this case only when both $B_{x}$ and $B_{z}$ are finite. It then leads to nonreciprocity in the critical supercurrents, $J_{c}^{+}\neq|J_{c}^{-}|$, hence exhibiting a finite diode efficiency ($\eta\neq 0$). Here, $J_{0}\equiv J_{c}(B_{z}=0)=|J_{c}^{+}(B_{z}=0)|=|J_{c}^{-}(B_{z}=0)|$. (b) and (c): The corresponding diode efficiency ($\eta$) is shown as a function of $B_{x}$ (with fixed $B_{z}$) and $B_{z}$ (with fixed $B_{x}$) with only linear Rashba SOC in the panels (b) and (c) respectively. The system parameters are chosen as: $(B_{y},\alpha,\mu/\Delta_{0},U,\beta^{-1})$=(00, $1$eV-Å, $0.5$, $25.55$ meV, $0.1$ meV).
@@
Figure 5: Superconducting diode effect with both linear and higher order SOC: (a) The supercurrent density ($J$) as a function of $q~{}$(nm-1) with both cubic Rashba SOC ($\delta$ = $\gamma$ = $100$ eV-Å3) and linear Rashba SOC ($\alpha=1$ eV-Å) choosing $B_{z}/\Delta_{0}=0$. The nonreciprocity in the critical supercurrents: $|J_{c}^{+}|\neq|J_{c}^{-}|$ (also implying a finite diode efficiency $\eta\neq 0$) is achieved for any finite value of $B_{x}$. Here, $J^{\prime}_{0}\equiv J_{c}(B_{x}=B_{z}=0)=|J_{c}^{+}(B_{x}=B_{z}=0)|=|J_{c}^{-}(B_{x}=B_{z}=0)|$. (b) and (c): The corresponding diode efficiency ($\eta$) is depicted as a function of $B_{x}$ (with fixed $B_{z}$) and $B_{z}$ (with fixed $B_{x}$) in the panels (b) and (c) respectively. We have chosen the other model parameters as: $(B_{y},\mu/\Delta_{0},U,\beta^{-1})$=(00, $0.5$, $16.45$ meV, $0.1$ meV).
@@ III Non-reciprocal transport
The supercurrent density $J(q)$ is computed from the condensation energy $\Omega(q,\Delta)$ in Eq. (6) by introducing a vector potential $A$ using minimal coupling and then taking the derivative of the condensation energy with respect to $A$ Yuan and Fu (2022) as,
$J(q)=-\frac{\partial\Omega(q+2eA,\Delta)}{\partial A}\bigg{|}_{A=0}=-2e\,\frac{\partial\Omega(q,\Delta)}{\partial q}.$ (9)
The sign of the $J(q)$ indicates the direction of the supercurrent flow. The critical values of the supercurrent along and opposite to the flow direction within the range $q_{c}^{-}\leq q\leq q_{c}^{+}$ for which the superconducting phase is stable, are denoted as $J_{c}^{+}$ and $J_{c}^{-}$ respectively. When $|J_{c}^{+}|\neq|J_{c}^{-}|$, we have a finite SDE in the system and the efficiency or the quality factor of the SDE is defined as
@@
We evaluate the diode efficiency ($\eta$) numerically, by first computing $J(q)$ over the range $q_{c}^{-}\leq q\leq q_{c}^{+}$ using the the self-consistency solutions $\Delta(q)$ for each value of $q$ shown in Fig 2(a) and Fig 3(a) for example, and then finding its optimal values $J_{c}^{+}$ and $J_{c}^{-}$. We systematically analyze the SDE in our Rashba nanowire device considering two broadly distinct scenarios: with only linear SOC and, with both linear and higher order SOC for various system parameters as discussed in the following two subsections.

### v2

@@ Abstract
The superconducting diode effect (SDE) refers to the non-reciprocal nature of the critical current (maximum current that a superconductor can withstand before turning into a normal metal) of a superconducting device. Here, we investigate SDE in helical superconductors with broken inversion and time-reversal symmetry, focusing on a prototypical Rashba nanowire device proximitized by an $s$-wave superconductor and subjected to external magnetic fields. Using a self-consistent Bogoliubov-de Gennes mean-field formalism, we analyze the interplay between linear and higher-order spin-orbit coupling (SOC), bulk supercurrents, and external magnetic fields. Our results demonstrate that Rashba nanowires with only linear SOC can achieve incredibly large diode efficiency $\gtrsim 45\%$ through the interplay of longitudinal and transverse magnetic fields. Notably, higher-order SOC enables finite diode efficiency even without a longitudinal Zeeman field, which can be utilized to reveal its presence and strength in nanowires. We present a comprehensive phase diagram of the device elucidating the emergent Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) superconducting state and demonstrate that proximitized Rashba nanowires offer a versatile, practical platform for SDE, with potential realizations in existing material systems. These results provide crucial insights for optimizing SDE in nanoscale superconducting devices, paving the way for next-generation dissipationless quantum electronics.
@@ I Introduction
The inversion symmetry breaking required for critical current non-reciprocity can be achieved through extrinsic mechanisms (e.g., asymmetric geometry or artificial superlattice potential Sundaresh et al. (2023); Burlakov et al. (2014); Narita and Ono (2024)) or intrinsic mechanisms like Rashba spin-orbit coupling (SOC) Ando et al. (2020); Yuan and Fu (2022); Ilić and Bergeret (2022); Turini et al. (2022); Roig et al. (2024); Halterman et al. (2022); Daido and Yanase (2022); Karabassov et al. (2022); Legg et al. (2022). Correspondingly, SDE can be broadly classified into two types: extrinsic and intrinsic. Although predicted decades ago Levitov et al. (1985); Edelstein (1996), the mechanism of SDE is beginning to be understood only recently, leading to extensive investigations into its potential for dissipationless electronic circuits Yuan and Fu (2022); Daido et al. (2022); Ilić and Bergeret (2022); He et al. (2022); Roig et al. (2024); Pal et al. (2022); de Picoli et al. (2023); Banerjee and Scheurer (2024); Nadeem et al. (2023). Recent theoretical studies Yuan and Fu (2022); Daido et al. (2022); Ilić and Bergeret (2022); He et al. (2022), using phenomenological Ginzburg-Landau theory and mean-field analysis of the two-dimensional Rashba-Zeeman Hubbard model, have shown that intrinsic SDE can arise in systems with broken inversion and time-reversal symmetry. In the helical superconducting ground state, this occurs due to asymmetry in Cooper pair momentum in the Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) state Fulde and Ferrell (1964); Larkin and Ovchinnikov (1965), resulting in Kramers non-degenerate $s$-wave Cooper pairs Daido et al. (2022); Yuan and Fu (2022); Nadeem et al. (2023).
@@
A Rashba nanowire device proximitized to an s-wave superconductor and subjected to external magnetic fields serves as a prototypical platform for realizing helical superconductivity Oreg et al. (2010); Sau et al. (2010); Lutchyn et al. (2010); Stanescu et al. (2011); Legg et al. (2022); de Picoli et al. (2023). While previous phenomenological/ non-self-consistent studies Legg et al. (2022) demonstrated the potential for SDE in such systems, they reported low diode efficiencies ($\sim 2\%$). To enhance the relatively low diode efficiency of this device and better understand its properties, it’s crucial to systematically investigate the superconducting order parameter using self-consistent methods Daido et al. (2022); Ilić and Bergeret (2022); de Picoli et al. (2023). Recent studies Daido et al. (2022); Ilić and Bergeret (2022); de Picoli et al. (2023) have shown that this approach can potentially uncover pathways for improvement across various parameters. In this paper, we address this gap through a comprehensive, self-consistent mean-field investigation of the interplay between various Rashba SOC terms and external magnetic fields in determining emergent FFLO pairing and highly efficient superconducting diodes in a Rashba nanowire device. We show that even with only linear Rashba SOC we can achieve a very large diode efficiency ($\gtrsim 45\%$) by suitably tuning the parameters in this system. We extend beyond linear Rashba SOC to include higher-order (cubic) Rashba SOC terms, which are not only naturally present in semiconducting nanowires Campos et al. (2018); Kammermeier et al. (2018); Sonehara et al. (2022) but also can dominate the normal state properties Vajna et al. (2012); Moriya et al. (2014); Nakamura et al. (2012); Michiardi et al. (2015); Shanavas (2016); Liu et al. (2018); Peng et al. (2023); Amundsen et al. (2024). Notably, we demonstrate that higher-order Rashba SOC can induce finite diode efficiency ($\gtrsim 15\%$) even without longitudinal magnetic fields. This behavior differs qualitatively from the linear SOC case and can be used to detect the presence of higher-order SOC terms in the nanowire. Our study establishes the Rashba nanowire device as a promising platform for highly efficient superconducting diodes, crucial for advancing nanoscale superconducting electronics.
@@ II Model and Self-consistent mean-field formalism
We consider a single channel one-dimensional semiconducting nanowire along the $z$-direction with both linear and higher order Rashba SOC terms Campos et al. (2018); Kammermeier et al. (2018); Sonehara et al. (2022); hso placed in close proximity to a three dimensional bulk $s$-wave superconductor as schematically shown in Fig. 1(a). In the presence of externally applied magnetic fields, the normal state physics of the nanowire is governed by the one-dimensional Rashba-Zeeman Hamiltonian (Alicea, 2012)
@@
$\!\!\hat{h}(z)=\frac{p_{z}^{2}}{2m}-\mu+\left(\frac{\alpha}{\hbar}p_{z}+\frac{\delta}{\hbar^{3}}p_{z}^{3}\right)\sigma_{x}+\frac{\gamma}{\hbar^{3}}p_{z}^{3}\sigma_{y}+\mathbf{B}\cdot\boldsymbol{\sigma}\,.$ (1)
Here, the spin degree of freedom is labeled by $s,s^{\prime}\in\{\uparrow,\downarrow\}$. $\psi_{s}^{\dagger}(z)$ creates a fermion at the position $z$ with spin $s$, while $\boldsymbol{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of the Pauli matrices in spin space and $\mu$ is the chemical potential. $p_{z}=\frac{\hbar}{i}\partial_{z}$ is the momentum operator in the direction of current flow ($z$-direction) within the nanowire. The strength of the usual linear Rashba SOC is denoted by $\alpha$ and the cubic Rashba SOC of strengths are $\delta$ and $\gamma$. The presence of Rashba SOC breaks the inversion symmetry of the system. The externally applied magnetic fields lead to the Zeeman term characterized by $\mathbf{B}=(B_{x},B_{y},B_{z})$ that breaks time-reversal symmetry. We assume the radius of the nanowire to be small such that there is only one occupied mode, i.e.,  single channel. This nanowire Hamiltonian [Eq. (1)] in momentum space can be written as
@@
$\hat{h}(k)=\xi_{k}+\left[\left(\alpha k+\delta k^{3}\right)\sigma_{x}+\gamma k^{3}\sigma_{y}\right]+\mathbf{B}\cdot\boldsymbol{\sigma}\,.$ (2)
@@
where, $\pm$ label the two helicty bands. Note that, the Zeeman field $B_{z}$ (longitudinal field) opens a Zeeman gap while $B_{x}$ (transverse field) creates asymmetry between the two bands as schematically shown in Fig. 1(b). On the other hand, $B_{y}$ opens up a gap as well as creates band asymmetry (not shown). Throughout the majority of this paper, we will discuss the results with $B_{y}=0$ for simplicity and will separately discuss the effects of nonzero $B_{y}$ for completeness.
We seek to elucidate the nature of the superconducting ground state in the Rashba nanowire in the presence of $s$-wave superconducting correlations that can be induced via the proximity effect in presence of a bulk $s$-wave superconductor, as well as supercurrents (see Fig. 1(a)). In general, the presence of supercurrents can lead to a superconducting state with finite momentum Cooper pairs termed as the FFLO state. To establish that, we consider an attractive Hubbard type interaction in the bulk $s$-wave superconductor described by the Hamiltonian
@@
Figure 2: FFLO ground state with linear SOC: (a) The self-consistent superconducting gap $\Delta(q)$ is shown as a function of the Cooper-pair momentum $q~{}$(nm-1) for $B_{x}/\Delta_{0}=0.3$ with only linear Rashba SOC ($\alpha=100$ meV-nm). $\Delta_{0}$ is the BCS gap in the absence of any external magnetic field ($\mathbf{B}=0$). (b) The Cooper pair momentum of the FFLO ground state $q_{0}$ as a function of $B_{x}$. The presence of the FFLO ground state ($q_{0}\neq 0$) is established only when both $B_{x}$ and $B_{z}$ are finite. Other model parameters are chosen as: $(B_{y},\mu/\Delta_{0},U,\beta^{-1})$=(00, $0.5$, $25.55$ meV, $0.1$ meV).
Figure 3: FFLO ground state with both linear and higher order SOC: (a) The self-consistent superconducting gap $\Delta(q)$ is depicted as a function of Cooper-pair momentum $q~{}$ for $B_{z}=0$ with both the linear Rashba SOC term ($\alpha=100$ meV-nm) and the cubic Rashba SOC terms ($\delta=\gamma=10$ eV-nm3). (b) The Cooper pair momentum of the FFLO ground state $q_{0}~{}$ as a function of $B_{x}$ is shown. Note that the FFLO state is stable even when $B_{z}=0$ in the presence of a finite $B_{x}$. Other system parameters are chosen as: $(B_{y},\mu/\Delta_{0},U,\beta^{-1})$=(00, $0.5$, $16.45$ meV, $0.1$ meV).
@@
Figure 4: Superconducting diode effect with linear SOC: (a) The supercurrent density ($J$) is shown as a function of $q~{}$(nm-1) considering solely linear Rashba SOC and $B_{x}/\Delta_{0}=0.3$. The nonreciprocity in the supercurrent, $J(q)\neq-J(-q)$, is achieved in this case only when both $B_{x}$ and $B_{z}$ are finite. It then leads to nonreciprocity in the critical supercurrents, $J_{c}^{+}\neq|J_{c}^{-}|$, hence exhibiting a finite diode efficiency ($\eta\neq 0$). Here, $J_{0}\equiv J_{c}(B_{z}=0)=|J_{c}^{+}(B_{z}=0)|=|J_{c}^{-}(B_{z}=0)|$. (b) and (c): The corresponding diode efficiency ($\eta$) is shown as a function of $B_{x}$ (with fixed $B_{z}$) and $B_{z}$ (with fixed $B_{x}$) with only linear Rashba SOC in the panels (b) and (c) respectively. The system parameters are chosen as: $(B_{y},\alpha,\mu/\Delta_{0},U,\beta^{-1})$=(00, $100$ meV-nm, $0.5$, $25.55$ meV, $0.1$ meV).
@@
Figure 5: Superconducting diode effect with both linear and higher order SOC: (a) The supercurrent density ($J$) as a function of $q~{}$(nm-1) with both cubic Rashba SOC ($\delta$ = $\gamma$ = $10$ eV-nm3) and linear Rashba SOC ($\alpha=100$ meV-nm) choosing $B_{z}=0$. The nonreciprocity in the critical supercurrents: $|J_{c}^{+}|\neq|J_{c}^{-}|$ (also implying a finite diode efficiency $\eta\neq 0$) is achieved for any finite value of $B_{x}$. Here, $J^{\prime}_{0}\equiv J_{c}(B_{x}=B_{z}=0)=|J_{c}^{+}(B_{x}=B_{z}=0)|=|J_{c}^{-}(B_{x}=B_{z}=0)|$. (b) and (c): The corresponding diode efficiency ($\eta$) is depicted as a function of $B_{x}$ (with fixed $B_{z}$) and $B_{z}$ (with fixed $B_{x}$) in the panels (b) and (c) respectively. We have chosen the other model parameters as: $(B_{y},\mu/\Delta_{0},U,\beta^{-1})$=(00, $0.5$, $16.45$ meV, $0.1$ meV).
@@ III Non-reciprocal transport
The supercurrent density $J(q)$ is computed from the condensation energy $\Omega(q,\Delta)$ in Eq. (6) as Yuan and Fu (2022); Daido et al. (2022); Legg et al. (2022),
$J(q)=-2e\,\frac{\partial\Omega(q,\Delta)}{\partial q}.$ (9)
The sign of $J(q)$ indicates the direction of the supercurrent flow. The critical (optimal) values of the supercurrent along and opposite to the flow direction within the range $q_{c}^{-}\leq q\leq q_{c}^{+}$ for which the superconducting phase is stable, are denoted as $J_{c}^{+}$ and $J_{c}^{-}$ respectively. When $|J_{c}^{+}|\neq|J_{c}^{-}|$, we have a finite SDE in the system and the efficiency or the quality factor of the SDE is defined as
@@
We first compute $J(q)$ over the range $q_{c}^{-}\leq q\leq q_{c}^{+}$ using the self-consistent solutions $\Delta(q)$ for each value of $q$ (shown in Fig 2(a) and Fig 3(a) for example). Then finding its optimal values $J_{c}^{+}$ and $J_{c}^{-}$, we evaluate the diode efficiency ($\eta$) numerically. We systematically analyze the SDE in our Rashba nanowire device considering two broadly distinct scenarios: with only linear SOC and, with both linear and higher order SOC for various system parameters as discussed in the following two subsections.